OpenStudy (anonymous):

FOILing, do i multiply or add? Like (2x+4)+(2x+4) i add there, right? And (2x+4)*(2x+4) I multiply there, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I did that in class todya but i didnt get the right answer..

OpenStudy (anonymous):

what did you get for the first one?

OpenStudy (anonymous):

(2x+4)+(2x+4) 4x+2x+4+4+2x+4+4 8x+16

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

it looks like you are multiplying and adding instead of adding. the parentheses here are unnecessary

OpenStudy (amistre64):

when it says to add, I would add. and when it says to multiply, id multiply. but thats just me set in my ways :)

OpenStudy (anonymous):

\[(2x+4)+(2x+4)=2x+4+2x+4=2x+2x+4+4\]

OpenStudy (anonymous):

let's imagine for a moment that x was ten (x is a variable, but it could be ten) then \[2x+4=2\times 10+4=24\] and \[24+24=48\]

OpenStudy (anonymous):

I thought it was like this